OpenStudy (anonymous):

It takes Trevon ten hours to clean an attic. Cody can clean the same attic in seven hours. Find how long it would take them if they worked together

OpenStudy (anonymous):

Trevon takes 10 hours, so every hour he cleans 1/10th of the attic (or 0.1 out of 1) Similarly Codey cleans 1/7th of the attic per hour (0.14286ish out of 1) So if they are both cleaning, the attic is being cleaned at a rate of 1/10th + 1/7th per hour = 0.1+0.14286=0.24286 per hour. We want to find how many hours it takes until the attic is cleaned. Call t the number of hours, then after t hours, t*0.24286 of the attic has been cleaned. So t*0.24846=1, and so t=1/(0.24846)=4.118 = 4 hours and 0.118 of an hour = 4 hours and 7 minutes. Sorry if this isn't very clear... I'm not great at explaining things :")